# Sylvester-Preconditioned Adaptive-Rank Implicit Time Integrators for Advection-Diffusion Equations with Inhomogeneous Coefficients

### v1

@@ 3.4 Truncation, Efficient Residual Computation, and Algorithm Flowchart
$=\underbrace{\left\|R_{U}\begin{bmatrix}{S_{1}}-\tilde{B}_{1}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&-\Delta t\,{S_{1}}&\mathbf{0}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&-\Delta t\,{S_{1}}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&-\Delta t\,{S_{1}}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&-\Delta t\,{S_{1}}\\ \end{bmatrix}R_{V}^{\top}\right\|}_{\in\mathbb{R}^{5r_{1}\times 5r_{1}}}$
$=\underbrace{\left\|R_{U}\begin{bmatrix}{S_{1}}-\tilde{B}_{1}&\mathbf{0}\\ \mathbf{0}&-\Delta t\,(I_{\mathcal{R}_{\text{ranks}}}\otimes S_{1})\\ \end{bmatrix}R_{V}^{\top}\right\|}\leq\epsilon_{\texttt{tol}},$
where $\mathcal{R}_{\text{ranks}}={\ell_{x}+k_{x}+\ell_{y}+k_{y}}$ is the total rank number of the advection and diffusion coefficients, and where $R_{U}$ and $R_{V}$ are the upper triangular matrices arising from the following QR decomposition:
{_$\{\_\ ,R_{U}\}$ $=\texttt{QR}\left(\begin{bmatrix}{U}_{1}&{T_{1}}{U}_{1}&\Phi_{1}{U}_{1}&{T_{3}}{U}_{1}&\Sigma_{1}{U}_{1}\end{bmatrix}\right),$ (23)
{_$\{\_\ ,R_{V}\}$ $=\texttt{QR}\left(\begin{bmatrix}{V}_{1}&\Phi_{2}{V}_{1}&{T_{2}}{V}_{1}&\Sigma_{2}{V}_{1}&{T_{4}}{V}_{1}\end{bmatrix}\right),$
where the orthogonal columns from the decomposition are ignored. The above residual result follows because, if ${\bf F}_{0}=U_{0}S_{0}V_{0}^{\top}$ (where $U_{0}\in\mathbb{R}^{N\times r_{0}}_{0}$ and $V_{0}\in\mathbb{R}^{N\times r_{0}}_{0}$ have orthonormal columns) then, since $U_{1}$ and $V_{1}$ contain $U_{0}$ and $V_{0}$, respectively, we have $U_{1}(U_{1})^{\top}{\bf F}_{0}V_{1}(V_{1})^{\top}={\bf F}_{0}$.
Input: Bases $U$, $V$; matrix of coefficients $S$; Tolerance $\epsilon$.
@@ 3.5 Generalization to Multi-rank Coefficients
We generalize next the adaptive-rank approach from rank-one advection and diffusion coefficients to the multi-rank case, using backward Euler. The corresponding pseudo-algorithm is presented in Algorithm 3. High-order temporal discretizations are considered later in this study. Discretization of the multi-rank advection-diffusion equation yields a multi-term generalized Sylvester equation of the form:
@@
$\alpha_{x,i}$ $=\text{average}(\Phi_{i}^{2,x}),\quad\alpha_{y,j}$ $=\text{average}(\Phi_{j}^{1,y}),\quad\gamma_{x,k}$ $=\text{average}(\Sigma_{k}^{2,x}),\quad\gamma_{y,l}$ $=\text{average}(\Sigma_{l}^{1,y}),$ (25)
to arrive at the approximated Sylvester equation:
@@
$\small P_{1}$ $=\sum_{i=1}^{\ell_{x}}{A_{1,i}}+\sum_{k=1}^{k_{x}}{A_{3,k}}=\sum_{i=1}^{\ell_{x}}{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{1}}-\Delta t\alpha_{x,i}T_{1,i}\right)}+\sum_{k=1}^{k_{x}}{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{1}}-\Delta t\gamma_{x,k}T_{3,k}\right)},$
$P_{2}$ $=\sum_{j=1}^{\ell_{y}}{A_{2,j}}+\sum_{l=1}^{k_{y}}{A_{4,l}}=\sum_{j=1}^{\ell_{y}}{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{2}}-\Delta t\alpha_{y,j}T_{2,j}\right)}+\sum_{l=1}^{k_{y}}{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{2}}-\Delta t\gamma_{y,l}T_{4,l}\right)}.$
The bases $U_{1}$ and $V_{1}$ are constructed from orthonormalization of dimension-wise generalized Krylov subspaces from equations (24) and (26) as follows:
$U_{1}$ $=\mathscr{T}_{\epsilon_{\kappa}}\left(\kappa_{m}\left(P_{1},P_{1}^{-1},A^{-1}_{1,1:\ell_{x}},A^{-1}_{3,1:k_{x}},\Phi^{1,y}_{1:\ell_{y}},\Sigma^{1,y}_{1:k_{y}},U_{0}\right)\right),$ (27)
$V_{1}$ $=\mathscr{T}_{\epsilon_{\kappa}}\left(\kappa_{m}\left(P_{2},P_{2}^{-1},A^{-1}_{2,1:\ell_{y}},A^{-1}_{4,1:k_{y}},\Phi^{2,x}_{1:\ell_{x}},\Sigma^{2,x}_{1:k_{x}},V_{0}\right)\right),$ (28)
In the multi-rank setting, after a Galerkin projection $U^{\top}\mathbf{R}_{\mathscr{{L}}}V=0$, we obtain a generalized Sylvester equation similar to (29) for the matrix of coefficients $S_{1}$:
${\bf S}_{1}-\Delta_{t}\left(\sum_{i=1}^{\ell_{x}}\tilde{T}_{1,i}{\bf S}_{1}\tilde{\Phi}_{i}^{2,x\top}+\sum_{j=1}^{\ell_{y}}\tilde{\Phi}_{j}^{1,y}{\bf S}_{1}\tilde{T}_{2,j}^{\top}+\sum_{k=1}^{k_{x}}\tilde{T}_{3,k}{\bf S}_{1}\tilde{\Sigma}_{k}^{2,x\top}+\sum_{l=1}^{k_{y}}\tilde{\Sigma}_{l}^{1,y}{\bf S}_{1}\tilde{T}_{4,l}^{\top}\right)=\tilde{{\bf B}}_{1}$ (29)
[…]
$\tilde{P}_{1}\mathbf{S}+\mathbf{S}\tilde{P}_{2}^{\top}=\tilde{B}_{1},$ (30)
@@
This yields the preconditioner operator:
$\mathscr{P}:\mathbf{S}\mapsto\tilde{P}_{1}\mathbf{S}+\mathbf{S}\tilde{P}_{2}^{\top}$ (31)
The whole process is summarized in Algorithm 3.
@@
Tolerances $\epsilon_{\kappa}$, $\epsilon$, $\epsilon_{\texttt{tol}}$;
Maximum iterations max_iter;
@@
1Compute operators $\{A_{1,i}\}_{i=1}^{\ell_{x}}$, $\{A_{2,j}\}_{j=1}^{\ell_{y}}$, $\{A_{3,k}\}_{k=1}^{k_{x}}$, $\{A_{4,l}\}_{l=1}^{k_{y}}$ according to Eq. (26);
2
3for $m=1$ to max_iter do
@@
5        Truncate to tolerance $\epsilon_{\kappa}$;
6        $U_{1}\leftarrow\mathscr{T}_{\epsilon_{\kappa}}\left(\kappa_{m}\left(P_{1},P_{1}^{-1},A^{-1}_{1,1:\ell_{x}},A^{-1}_{3,1:k_{x}},\Phi^{1,y}_{1:\ell_{y}},\Sigma^{1,y}_{1:k_{y}},U_{0}\right)\right)$;
7
8       $V_{1}\leftarrow\mathscr{T}_{\epsilon_{\kappa}}\left(\kappa_{m}\left(P_{2},P_{2}^{-1},A^{-1}_{2,1:\ell_{y}},A^{-1}_{4,1:k_{y}},\Phi^{2,x}_{1:\ell_{x}},\Sigma^{2,x}_{1:k_{x}},V_{0}\right)\right)$;
9
@@ 5        Truncate to tolerance $\epsilon_{\kappa}$;
11        Solve the reduced equation (29) for $S_{1}$ using GMRES-ACS (31) to tolerance $\epsilon_{\texttt{GMRES}}$;
@@
13        Truncate $\{U_{1},S_{1},V_{1}\}\leftarrow\mathcal{T}_{\epsilon}(\{U_{1},S_{1},V_{1}\})$ to tolerance $\epsilon$;
@@
15        Compute {_$\{\_,R_{U}\}=\texttt{QR}\left([U_{1},T_{1,1:\ell_{x}}U_{1},\Phi_{1:\ell_{y}}^{1,y}U_{1},T_{3,1:k_{x}}U_{1},\Sigma_{1:k_{y}}^{1,y}U_{1}]\right)$ and
16        {_$\{\_,R_{V}\}=\texttt{QR}\left([V_{1},\Phi_{1:\ell_{x}}^{2,x}V_{1},T_{2,1:\ell_{y}}V_{1},\Sigma_{1:k_{x}}^{2,x}V_{1},T_{4,1:k_{y}}V_{1}]\right)$ ;
17
18       Compute the residual $\|\mathbf{R}\|=\left\|R_{U}\text{diag}\left(S_{1}-\tilde{B}_{1},-\Delta t\left(I_{\mathcal{R}_{\text{ranks}}}\otimes S_{1}\right)\right)R_{V}^{\top}\right\|$;
@@
20               Reject solution and return to augment bases further;
21
@@ 25               Exit loop;
26
27
Algorithm 3 Backward Euler Adaptive-Rank Integrator for the Generalized Sylvester Equation. 00footnotetext: The MATLAB implementation of the algorithm is publicly available on GitHub, Adaptive-rank-ACS, https://github.com/helkahza1/Adaptive-rank-ACS
@@ 3.6 Computational Complexity
We now analyze the computational complexity of Algorithm 3 for backward Euler. For conciseness, we consider the same spatial resolution per dimension, i.e., $N=N_{1}=N_{2}$. Starting from an initial condition $U_{0}S_{0}V_{0}^{\top}$, where $U_{0}\in\mathbb{R}^{N\times r_{0}}$, $V_{0}\in\mathbb{R}^{N\times r_{0}}$, and $S_{0}\in\mathbb{R}^{r_{0}\times r_{0}}$, the $m$-th extended Krylov iteration generates $U_{1}\in\mathbb{R}^{N\times r_{\kappa}}$, $V_{1}\in\mathbb{R}^{N\times r_{\kappa}}$, and $S_{1}\in\mathbb{R}^{r_{\kappa}\times r_{\kappa}}$, where $r_{\kappa}=((2+\mathcal{R}_{\text{ranks}})m+1)r_{0}$. Using SVD-truncated QR, this dimension is then reduced to smaller ranks, $r_{x}$ and $r_{y}$, for the $x$ and $y$ directions, respectively. The matrix $S_{1}$ from the $m$-th Krylov iteration solve, initially of size $r_{x}\times r_{y}$, is truncated to rank $r_{1}\leq\min(r_{x},r_{y})$ using Algorithm 2. To simplify the complexity analysis, similar to Section 3.3, let $r=\max(r_{x},r_{y})$. The computational complexity of each step is then estimated as follows:
@@
Constructing the generalized Krylov basis as outlined in lines (3) and (3) of Algorithm 3 requires performing the operations $A_{1,1:r_{x}}^{-1}U_{0}$ and $A_{3,1:k_{x}}^{-1}U_{0}$, and $A_{2,1:r_{y}}^{-1}V_{0}$ and $A_{4,1:k_{y}}^{-1}V_{0}$ which necessitates solving the systems $AX=U$ and $AX=V$, where the subscripts have been omitted for simplicity. This is typically achieved through an LU factorization of the operator. With 1D finite differences, $A$ has a tridiagonal structure, which can be efficiently factorized using techniques such as the Thomas algorithm, resulting in a complexity of $\mathcal{O}(N)$ per column of $X$ [28]. The same complexity holds for applying $P_{1}^{-1}$ and $P_{2}^{-1}$. Additionally, applying $P_{1}U_{0}$, $\Phi_{1,3}U_{0}$, $\Sigma_{1,3}U_{0}$, $P_{2}V_{0}$, $\Phi_{2,4}V_{0}$, and $\Sigma_{2,4}V_{0}$, can be done with $\mathcal{O}(Nr_{0})$ computational complexity for sparse diagonal and tridiagonal matrices, leading to a total complexity of $\mathcal{O}(N\mathcal{R}_{\text{ranks}}r_{0})$ to form the extended Krylov subspaces.
Regarding truncated orthonormalization in Algorithm 1, using a modified Gram-Schmidt method to compute the reduced QR decomposition of $U_{1}$ and $V_{1}$, in line (1), requires a complexity of $\mathcal{O}\left(Nr_{\kappa}^{2}\right)$. The SVD decomposition of the small-sized upper triangular matrix in line (1) requires complexity $\mathcal{O}\left(r_{\kappa}^{3}\right)$. The basis update in line (1) requires a complexity of $\mathcal{O}\left(Nr_{\kappa}r\right)$, where ${{r}}\leq r_{\kappa}$. Hence, the complexities of lines (3), (3) in Algorithm 3 scale as $\mathcal{O}\left(Nr_{\kappa}^{2}\right)$. Generally, we expect $r_{\kappa}\ll N$ when the solution exhibits a low-rank structure.
@@
The application of the operator using matrix-matrix multiplication and summation in the Krylov method (GMRES), as outlined in Section, 3.3, requires a complexity of $\mathcal{O}(r^{3})$. Furthermore, applying the preconditioner involves solving a Sylvester equation of size $r$, which can be achieved using the Bartels-Stewart algorithm [11] with a complexity of $\mathcal{O}(r^{3})$. Hence, the $S$-step has $\mathcal{O}(r^{3})$ complexity.
@@
The SVD decomposition of ${\widetilde{S}}$ in line (2) of Algorithm 2 is of $\mathcal{O}(r^{3})$, also independent of $N$. Updating the bases in lines (2) and (2) of Algorithm 2 requires $\mathcal{O}(Nr{r_{1}})$, where ${r_{1}}\leq r$ is the post-truncation rank of the new solution as described in Algorithm 2.

### v2

@@ 3.4 Truncation, Efficient Residual Computation
$=\underbrace{\left\|R_{U}\begin{bmatrix}-\tilde{B}_{1}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&{{S}_{1,\epsilon}}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&-\Delta t\,{{S}_{1,\epsilon}}&\mathbf{0}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&-\Delta t\,{{S}_{1,\epsilon}}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&-\Delta t\,{{S}_{1,\epsilon}}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&-\Delta t\,{{S}_{1,\epsilon}}\\ \end{bmatrix}R_{V}^{\top}\right\|}_{\in\mathbb{R}^{(r_{x}+5r_{1})\times(r_{y}+5r_{1})}}$
$={\left\|R_{U}\begin{bmatrix}-\tilde{B}_{1}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&{S_{1,\epsilon}}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&-\Delta t\,(I_{\mathcal{R}_{\text{ranks}}}\otimes{S_{1,\epsilon}})\end{bmatrix}R_{V}^{\top}\right\|}\leq\epsilon_{\texttt{tol}},$
where $\mathcal{R}_{\text{ranks}}={\ell_{x}+k_{x}+\ell_{y}+k_{y}}$ is the total rank number of the advection and diffusion coefficients, and $R_{U}$ and $R_{V}$ are the upper triangular matrices arising from the following Q-less QR decomposition:
{_$\{\_\ ,R_{U}\}$ $=\texttt{QR}\left(\begin{bmatrix}{U}_{1}&{U}_{1,\epsilon}&{T_{1}}{U}_{1,\epsilon}&\Phi_{1}{U}_{1,\epsilon}&{T_{3}}{U}_{1,\epsilon}&\Sigma_{1}{U}_{1,\epsilon}\end{bmatrix}\right),$ (24)
{_$\{\_\ ,R_{V}\}$ $=\texttt{QR}\left(\begin{bmatrix}{V}_{1}&{V}_{1,\epsilon}&\Phi_{2}{V}_{1,\epsilon}&{T_{2}}{V}_{1,\epsilon}&\Sigma_{2}{V}_{1,\epsilon}&{T_{4}}{V}_{1,\epsilon}\end{bmatrix}\right).$
The above residual result follows from the property that $U_{1}(U_{1})^{\top}{\bf F}_{0}V_{1}(V_{1})^{\top}={\bf F}_{0}$, with ${\bf F}_{0}=U_{0}S_{0}V_{0}^{\top}$ (where $U_{0}\in\mathbb{R}^{N\times r_{0}}_{0}$ and $V_{0}\in\mathbb{R}^{N\times r_{0}}_{0}$ have orthonormal columns), since $U_{1}$ and $V_{1}$ contain $U_{0}$ and $V_{0}$, respectively.
Input: Bases $U$, $V$; matrix of coefficients ${\bf S}$; Tolerance $\epsilon$.
@@ 3.5 Generalization to Multi-rank Coefficients
We generalize next the adaptive-rank approach to the multi-rank coefficient case in (5) and (6), using backward Euler for now. Discretization of the advection-diffusion equation in this case yields a multi-term GSE of the form:
@@
$\small P_{1}$ $=\sum_{i=1}^{\ell_{x}}{A_{1,i}}+\sum_{k=1}^{k_{x}}{A_{3,k}}=\sum_{i=1}^{\ell_{x}}{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{1}}-\Delta t\alpha_{x,i}T_{1,i}\right)}+\sum_{k=1}^{k_{x}}{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{1}}-\Delta t\gamma_{x,k}T_{3,k}\right)},$ (27)
$P_{2}$ $=\sum_{j=1}^{\ell_{y}}{A_{2,j}}+\sum_{l=1}^{k_{y}}{A_{4,l}}=\sum_{j=1}^{\ell_{y}}{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{2}}-\Delta t\alpha_{y,j}T_{2,j}\right)}+\sum_{l=1}^{k_{y}}{\left(\frac{1}{\mathcal{R}_{\text{ranks}}}I_{N_{2}}-\Delta t\gamma_{y,l}T_{4,l}\right)}.$ (28)
Also as before, the bases $U_{1}$ and $V_{1}$ are constructed from orthonormalization of dimension-wise xKrylov from equations (25) and (26) as:
$U_{1}$ $=\kappa^{\texttt{trunc}}_{m}\left(P_{1},P_{1}^{-1},A^{-1}_{1,1:\ell_{x}},A^{-1}_{3,1:k_{x}},\Phi^{1,y}_{1:\ell_{y}},\Sigma^{1,y}_{1:k_{y}},U_{0}\right),$ (29)
$V_{1}$ $=\kappa^{\texttt{trunc}}_{m}\left(P_{2},P_{2}^{-1},A^{-1}_{2,1:\ell_{y}},A^{-1}_{4,1:k_{y}},\Phi^{2,x}_{1:\ell_{x}},\Sigma^{2,x}_{1:k_{x}},V_{0}\right).$ (30)
In the multi-rank setting, after a Galerkin projection $U^{\top}\mathbf{R}_{\mathscr{{L}}}V=0$, we obtain a GSE similar to (21) for the matrix of coefficients ${\bf S}_{1}$:
${\bf S}_{1}-\Delta_{t}\left(\sum_{i=1}^{\ell_{x}}\tilde{T}_{1,i}{\bf S}_{1}\tilde{\Phi}_{i}^{2,x\top}+\sum_{j=1}^{\ell_{y}}\tilde{\Phi}_{j}^{1,y}{\bf S}_{1}\tilde{T}_{2,j}^{\top}+\sum_{k=1}^{k_{x}}\tilde{T}_{3,k}{\bf S}_{1}\tilde{\Sigma}_{k}^{2,x\top}+\sum_{l=1}^{k_{y}}\tilde{\Sigma}_{l}^{1,y}{\bf S}_{1}\tilde{T}_{4,l}^{\top}\right)=\tilde{{\bf B}}_{1},$ (31)
where the projected operators are analogous to those defined in the rank-one case. We solve (31) for the ${\bf S}_{1}$ matrix using preconditioned GMRES. The preconditioner is constructed by another Galerkin projection to find the projected SE:
$\tilde{P}_{1}\mathbf{S}+\mathbf{S}\tilde{P}_{2}^{\top}=\tilde{{\bf B}}_{1},$ (32)
@@
This yields the multi-rank-coefficients ACS preconditioner operator:
$\mathscr{P}:\mathbf{S}\mapsto\tilde{P}_{1}\mathbf{S}+\mathbf{S}\tilde{P}_{2}^{\top}.$ (33)
We summarize the pseudo-algorithm for multi-rank coefficients in Algorithm 4. The tolerances $\epsilon_{\kappa}$, $\epsilon_{\texttt{GMRES}}$, $\epsilon$, $\epsilon_{\texttt{tol}}$ in the algorithm are set as follows. First, we choose the residual tolerance for xKrylov outer iterations, $\epsilon_{\texttt{tol}}$. To prevent rapid growth of the xKrylov subspaces, we choose the Krylov basis truncation threshold as $\epsilon_{\kappa}\leq\epsilon_{\texttt{tol}}$, typically two to three orders of magnitude smaller. The GMRES-ACS tolerance for the inner reduced system solve, $\epsilon_{\texttt{GMRES}}$, is initially set to be less than $\epsilon_{\texttt{tol}}$ for the first outer-basis iteration. It is subsequently adjusted to be two to three orders of magnitude smaller than the relative residual error from the previous xKrylov outer iteration. This adjustment ensures that errors from the inner solve do not pollute the accuracy of the outer solve. The SVD truncation threshold $\epsilon$ is chosen to be one to three orders of magnitude smaller than the tolerance $\epsilon_{\text{tol}}$. Since solution truncation occurs before residual evaluations, the truncation error in the solution should remain below the residual tolerance to prevent residual pollution by over-truncation.
@@
Tolerances $\epsilon_{\kappa}$, $\epsilon_{\texttt{GMRES}}$, $\epsilon$, $\epsilon_{\texttt{tol}}$;
@@
1Compute operators $\{A_{1,i}\}_{i=1}^{\ell_{x}}$, $\{A_{2,j}\}_{j=1}^{\ell_{y}}$, $\{A_{3,k}\}_{k=1}^{k_{x}}$, $\{A_{4,l}\}_{l=1}^{k_{y}}$ according to Eqs. (27), (28);
2 Compute $l=2+\ell_{x}+\ell_{y}+k_{x}+k_{y}$;
3 Set $U^{(i)}=U_{0}$ and $V^{(i)}=V_{0}$ for $i=1,\dots,l$;
4 Set $U_{1}=U_{0}\quad V_{1}=V_{0}$;
5 while not converged do
@@ 4 Set $U_{1}=U_{0}\quad V_{1}=V_{0}$;
7        Orthogonalize and truncate to tolerance $\epsilon_{\kappa}$;
8        $U_{1},\{U^{(1)}\cdots U^{(l)}\}\leftarrow\kappa_{m}^{\texttt{trunc}}\left(\{P_{1},P_{1}^{-1},A^{-1}_{1,1:\ell_{x}},A^{-1}_{3,1:k_{x}},\Phi^{1,y}_{1:\ell_{y}},\Sigma^{1,y}_{1:k_{y}}\};\{U^{(1)}\cdots U^{(l)}\};U_{1};\epsilon_{\kappa}\right)$;
9        $V_{1},\{V^{(1)}\cdots V^{(l)}\}\leftarrow\kappa_{m}^{\texttt{trunc}}\left(\{P_{2},P_{2}^{-1},A^{-1}_{2,1:\ell_{y}},A^{-1}_{4,1:k_{y}},\Phi^{2,x}_{1:\ell_{x}},\Sigma^{2,x}_{1:k_{x}}\};\{V^{(1)}\cdots V^{(l)}\};V_{1};\epsilon_{\kappa}\right)$;
10
@@ 7        Orthogonalize and truncate to tolerance $\epsilon_{\kappa}$;
12        Solve the reduced GSE (31) for ${\bf S}_{1}$ using GMRES-ACS (33) to tolerance $\epsilon_{\texttt{GMRES}}$;
@@
14        Truncate $\{U_{1,\epsilon},S_{1,\epsilon},V_{1,\epsilon}\}\leftarrow\mathcal{T}_{\epsilon}(\{U_{1},S_{1},V_{1}\})$ to tolerance $\epsilon$ (Alg. 3);
15
@@
17        Compute {_$\{\_,R_{U}\}=\texttt{QR}\left([U_{1},U_{1,\epsilon},T_{1,1:\ell_{x}}U_{1,\epsilon},\Phi_{1:\ell_{y}}^{1,y}U_{1,\epsilon},T_{3,1:k_{x}}U_{1,\epsilon},\Sigma_{1:k_{y}}^{1,y}U_{1,\epsilon}]\right)$ and
18        {_$\{\_,R_{V}\}=\texttt{QR}\left([V_{1},V_{1,\epsilon},\Phi_{1:\ell_{x}}^{2,x}V_{1,\epsilon},T_{2,1:\ell_{y}}V_{1,\epsilon},\Sigma_{1:k_{x}}^{2,x}V_{1,\epsilon},T_{4,1:k_{y}}V_{1,\epsilon}]\right)$ ;
19
20       Compute the residual $\|\mathbf{R}\|=\left\|R_{U}\text{diag}\left(-\tilde{B}_{1},S_{1,\epsilon},-\Delta t\left(I_{\mathcal{R}_{\text{ranks}}}\otimes S_{1,\epsilon}\right)\right)R_{V}^{\top}\right\|$;
21
@@
23               Reject solution and cycle to augment bases further;
24
@@ 28               Exit loop;
29
30
Algorithm 4 Backward Euler Adaptive-Rank Integrator for the GSE.
@@ 3.6 Computational Complexity
We analyze next the computational complexity of Algorithm 4 for backward Euler. For conciseness, we consider the same spatial resolution per dimension, i.e., $N=N_{1}=N_{2}$. Starting from an initial condition $U_{0}S_{0}V_{0}^{\top}$, where $U_{0}\in\mathbb{R}^{N\times r_{0}}$, $V_{0}\in\mathbb{R}^{N\times r_{0}}$, and $S_{0}\in\mathbb{R}^{r_{0}\times r_{0}}$, the $m$-th xKrylov iteration generates $U_{1}\in\mathbb{R}^{N\times r_{\kappa}}$, $V_{1}\in\mathbb{R}^{N\times r_{\kappa}}$, and $S_{1}\in\mathbb{R}^{r_{\kappa}\times r_{\kappa}}$, where $r_{\kappa}=(2+\mathcal{R}_{\text{ranks}})r^{\ast}_{0}+r^{(m-1)}$. Here, $r^{(m-1)}$ denotes the basis size from the previous xKrylov iteration (initialized with $r^{(0)}=r_{0}$), and $r^{\ast}_{0}\leq r_{0}$ is the possibly truncated initial rank as produced by line (2) of Algorithm 2. Using SVD-truncated QR, this dimension is then reduced to smaller ranks, $r_{x}$ and $r_{y}$. For simplicity, let $\tilde{r}=\max(r_{x},r_{y})$. The matrix $S_{1}$ from the $m$-th inner iteration solve, initially of size $r_{x}\times r_{y}$, is truncated to rank $r_{1}\leq\min(r_{x},r_{y})$ using Algorithm 3. To simplify the complexity analysis, let $r=\max(r_{0},r^{(m-1)},\tilde{r},(\mathcal{R}_{\text{ranks}}+1)r_{1})$, which we consider independent of the mesh-refinement $N$ for practical purposes. The computational complexity of each step is then estimated as follows:
@@
Constructing the generalized Krylov basis as outlined in lines (4) and (4) of Algorithm 4 requires performing the operations $A_{1,1:r_{x}}^{-1}U_{0}$ and $A_{3,1:k_{x}}^{-1}U_{0}$, and $A_{2,1:r_{y}}^{-1}V_{0}$ and $A_{4,1:k_{y}}^{-1}V_{0}$, which require solving the systems $AX=U$ and $AX=V$ (subscripts have been omitted for simplicity). With 1D finite-differences, $A$ has a tridiagonal structure, which can be efficiently factorized using direct-solver techniques such as the Thomas algorithm, with $\mathcal{O}(N)$ complexity per column of $X$ [34]. The same complexity holds for applying $P_{1}^{-1}$ and $P_{2}^{-1}$. Additionally, applying $P_{1}U_{0}$, $\Phi_{1,3}U_{0}$, $\Sigma_{1,3}U_{0}$, $P_{2}V_{0}$, $\Phi_{2,4}V_{0}$, and $\Sigma_{2,4}V_{0}$, can be done with $\mathcal{O}(Nr^{\ast}_{0})$ computational complexity for sparse diagonal and tridiagonal matrices, leading to a total complexity of $\mathcal{O}(N(2+\mathcal{R}_{\text{ranks}})r^{\ast}_{0})$ to augment the xKrylov subspaces. The use of Algorithm 2 to compute the orthonormal truncated bases $U_{1}$ and $V_{1}$ incurs a complexity of $\mathcal{O}\left(N\,\max\left[\left(2+\mathcal{R}_{\text{ranks}}\right)\,r_{0}^{\ast}\,r^{(m-1)},\;\left(\left(2+\mathcal{R}_{\text{ranks}}\right)\,r_{0}^{\ast}\right)^{2}\right]\right)\sim\mathcal{O}(Nr^{2})$. Generally, we expect $r$ to be much smaller and independent of $N$ when the solution is sufficiently smooth and exhibits a low-rank structure.
@@
The application of the reduced GSE operator using matrix-matrix multiplication and summation in the Krylov method (GMRES) features a complexity of $\mathcal{O}\left(\tilde{r}^{d+1}\right)$. Applying the ACS preconditioner requires solving an SE of size $r_{x}\times r_{y}$, which can be achieved using the Bartels-Stewart algorithm [15] with a complexity of $\mathcal{O}(\tilde{r}^{d+1})$. Hence, the $S$-step has $\mathcal{O}(r^{d+1})$ complexity.
@@
The SVD decomposition of ${\widetilde{S}}$ in line (3) of Algorithm 3 is of $\mathcal{O}(\tilde{r}^{d+1})$ complexity. Updating the bases in lines (3) and (3) of Algorithm 3 requires $\mathcal{O}(N\tilde{r}{r_{1}})$ operations, where ${r_{1}}\leq\min(r_{x},r_{y})$ is the post-truncation rank of the new solution from Algorithm 3.